OpenStudy (anonymous):

divide 6/7x slash 1/3x? plz explain.

OpenStudy (lgbasallote):

\[\huge \frac{\frac{6}{7x}}{\frac{1}{3x}} \implies \frac{6}{7x}\;\times\; \frac{3x}{1}\] was that clear?

OpenStudy (anonymous):

kind of, and the answer was not 1 of my choices?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

I'll show you.

OpenStudy (lgbasallote):

sue

OpenStudy (lgbasallote):

sure*

OpenStudy (anonymous):

25/21x , 2/7x , 11/21x and 5/4x are my choices I want to understand how to work it out and the answer you gave was not one of these. srry.

OpenStudy (lgbasallote):

well did the question i put correct

OpenStudy (lgbasallote):

i may have misunderstood your question

OpenStudy (anonymous):

do you want me to draw it?

OpenStudy (lgbasallote):

by all means

OpenStudy (anonymous):